OpenStudy (anonymous):

PLEASE HELP!!! let f(x)=(x^2+3x-4) and g(x)=(x+4) a. find f*g and state the domain b. find f/g and state the domain

OpenStudy (tkhunny):

Well, do it. What is f*g?

OpenStudy (anonymous):

Thats just it im not sure

OpenStudy (tkhunny):

I do have to wonder why you have been given such a problem statement with NO introduction to the notation. That just makes no sense. f*g = f(x) * g(x) Now, can you produce f*g?

OpenStudy (anonymous):

alright but how do i do f(x) when im not given an x to begin with

OpenStudy (tkhunny):

f/g = f(x) / g(x) = (x^2+3x-4) / (x+4)

OpenStudy (tkhunny):

You are given a definition of f(x). You do NOT need a specific value for 'x'. The definition of f(x) is a relation for ALL values of x. You are being asked to consider and algebraic manipulation, not a numerical result.

OpenStudy (anonymous):

alright

OpenStudy (tkhunny):

Your turn. What is f*g?

OpenStudy (anonymous):

x^3+7x^2+8x-16??

OpenStudy (tkhunny):

Very good. It may have been sufficient to write f*g = (x^2+3x-4)*(x+4). Okay, that is your, basic, run-of-the-mill polynomial Relation. What is the Domain of such a relation?

OpenStudy (anonymous):

wouldnt that be all real numbers?

OpenStudy (tkhunny):

Excellent. All Real Numbers. How about f/g?

OpenStudy (anonymous):

x-1??

OpenStudy (tkhunny):

No, that's no good. It may be closely related to h(x) = x-1, but it is not quite the same thing. Please do the following exercise: If h(x) = x - 1 And if r(x) = (x^2+3x-4) / (x+4) Find h(-4) and r(-4). Let's see if we learning anything that is not obvious at the moment.

OpenStudy (anonymous):

h(-4)=-5 and 0 right?

OpenStudy (tkhunny):

Nope. Try r(-4) again. Do the numerator and denominator separately. Perhaps you will see the problem.

OpenStudy (tkhunny):

That was pretty good, though. We did establish that h(x) and r(x) are NOT the same thing. You just told be -5 and 0. Not the same.

OpenStudy (anonymous):

i just did that and got that same thing

OpenStudy (anonymous):

it doesnt seem to work

OpenStudy (tkhunny):

You should not get zero. You are not understanding what you are seeing. Let's learn something about zero (0). 12/4 = 3 can be expressed as 3*4 = 12 6/3 = 2 can be expressed as 2*3 = 6 14/2 = 7 can be expressed as 2*7 = 14 -12/2 = -6 can be expressed as -6(2) = -12 Make sense so far? Bear with me. It will be worth it when we get there!

OpenStudy (anonymous):

okay

OpenStudy (tkhunny):

Is there another expression for 0/0?

OpenStudy (anonymous):

0*0??

OpenStudy (tkhunny):

If 0/0 = 0, we should be able to express it as 0*0 = 0, but isn't it also the case that we can express it as 0*5 = 0 or 0*12 = 0 or 0*(-6) = 0? It doesn't really matter what that other value is. If the first is zero, there is no significance to the second value. In the second value, 0 just isn't important. Any sense at all, yet?

OpenStudy (anonymous):

yes

OpenStudy (tkhunny):

Here's the kicker. Is is very important that you get this one stuck in your head. Zero in the denominator doesn't mean anything. It just doesn't matter what's int eh numerator. 0/0 is NOT 1. It's also not "infinity". It's also not 5 and it's not the square root of 2. When you evaluated r(-4), you should get zero for both numerator and denominator separately. I believe you did that, but then you wanted to interpret that as 0/0 = 0 and that is no good. What we have is this: f/g is a rational function with x = -4 NOT in the Domain. You absolutely may not use x = -4 in f/g. When you wrote f/g = x-1 you were ALMOST correct. This is correct for EVERY Real Number EXCEPT x = -4. As long as you avoid x = -4, then f/g = x-1. Soaking it all up? It is glorious to behold!

OpenStudy (anonymous):

so then the domain is \[x \neq 4\] and f/g=x-1??

OpenStudy (anonymous):

all real numbers except -4